OpenStudy (anonymous):

Help finding the names of polygons please if i am given the interior or exterior how can i know the name of the polygon can someone give me a formula or website please

ganeshie8 (ganeshie8):

if exterior angle of regular polygon is \(\large \mathbb{x}\) degrees, then : number of sides in polygon = \(\large \mathbb{\frac{360}{x}}\)

ganeshie8 (ganeshie8):

you need to remember oly that

OpenStudy (anonymous):

wait is it \[x = \frac{ 360 }{ x }\] no matter how big or small the number is @ganeshie8 ?

ganeshie8 (ganeshie8):

yes, number of sides = \( \huge \frac{360}{exterior ~~angle}\)

OpenStudy (anonymous):

okay if i am given exterior angle is 24 than i would divide by 24 and get 15 sides than what equation would i use to find the interior angles? ( 15 - 2) 180 one?

ganeshie8 (ganeshie8):

use this : interior angle + exterior angle = 180

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

their sum adds up to 180

ganeshie8 (ganeshie8):

so if u knw exterior angel, simply subtract it from 180, u wud get interior angle !

OpenStudy (anonymous):

Thank you very much! i get it now great help

ganeshie8 (ganeshie8):

glad to hear :)